i was given the functions of f(x)=3x-5 and g(x)=1/3 - 4/3x. i have to correct the errors in this problem. g(-3)=1/3 - 4/3(3)
=1/3 -4
=-3 and 2/3
where is it wrong

Answer by Alan3354(69443)

f(x)=3x-5 and g(x)=1/3 - 4/3x
----------
Is it g(x) = 1/3 - (4/3)x ?
or g(x) = 1/3 - 4/(3x) ?
Parentheses are free
-------------
Using
g(x) = 1/3 - (4/3)x
g(-3) = 1/3 - (4/3)*(-3)
= 1/3 + 4
= 13/3